OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

K to find we can solve this problem by the using elimination. We can multiply the first equation to make the y's cancel out. -2(2x-y=6)= -4x+2y=-12 4x-2y=12 Notice when we try to subtract this, it comes out to 0=0. This means that there is an infinite amount of solutions so your answer would be C-(x can be any number as there are infinitely many solutions to the system.)

OpenStudy (anonymous):

Thank You so much! ^^

OpenStudy (anonymous):

No problem :)